14. One card is drawn from a well-shuffled deck of 52 cards. Find the probability of getting

(i) a king of red colour

(ii) a face card

(iii) a red face card

(iv) the jack of hearts

(v) a spade

(vi) the queen of diamonds

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Total number of possible outcomes = 52 P(E) = (Number of favourable outcomes/ Total number of outcomes) (i) Total numbers of king of red colour = 2 P (getting a king of red colour) = 2/52 = 1/26 = 0.038 (ii) Total numbers of face cards = 12 P (getting a face card) = 12/52 = 3/13 = 0.23 (iii) Total numbers of red face cards = 6 P (getting a king of red colour) = 6/52 = 3/26 = 0.11 (iv) Total numbers of jack of hearts = 1 P (getting a king of red colour) = 1/52 = 0.019 (v) Total numbers of king of spade = 13 P (getting a king of red colour) = 13/52 = ¼ = 0.25 (vi) Total numbers of queen of diamonds = 1 P (getting a king of red colour) = 1/52 = 0.019